OpenStudy (anonymous):

Find all the zeros of the equation. -4x4 - 44x2 + 3600 = 0 A. 5, -5, 6i, -6i B. 5, 6i C. 5, -5, 6i, 0 D. -5, -6i

OpenStudy (anonymous):

4x^4+44x^2-3600=0 x^4+11x^2-900=0 x^4+36x^2-25x^2-900=0 x^2(x^2+36)-25(x^2+36)=0 (x^2-25)(x^2+36)=0 so x^2=25,-36 x=5,-5,6i,-6i

OpenStudy (anonymous):

Yeah i thought it was a :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

you are welcome!